solve by any method you prefer.
3w^2 - 16 = 0

Answer by fractalier(2101)

From
3w^2 - 16 = 0
add 16, divide by three and take the root and get
3w^2 = 16
w^2 = 16/3
w = ± sqrt(16/3) now rationalize
w = ± 4√3 / 3